OpenStudy (anonymous):

Given the following, find the function of g: f(x) = 1 (f ) x+7 -- __ (x) = ----- x g x^2-x g(x) = ?

OpenStudy (anonymous):

g(x) = x- 1 / x +7

OpenStudy (anonymous):

can you show me how you got that?

OpenStudy (anonymous):

divide both sides by f(x) to get 1 / g(x) = x+7/ x -1 now just take the reciprocal

OpenStudy (anonymous):

u r welcome :D

jhonyy9 (jhonyy9):

f(x)=1/x f(x) x+7 x+7 --- = ------ = ------ so if we sign g(x) with y we get g(x) x*2 - x x(x-1) 1/x x+7 X-1 --- = ------ ----- (x(x-1))/X = Y(X+7) ----- X-1= Y(X+7) ---- Y= ---- y x(x-1) X+7 x-1 g(x)= ----- x+7